Thomas’ Calculus Early Transcendentals 13ed

Homework Assignment for Chapter 15

1. Section 15.1: Problems 7, 21.

2. Section 15.2: Problems 11, 19, 27, 29, 35, 43, 47, 69.

3. Section 15.3: Problems 13, 17.

4. Section 15.4: Problems 11, 21, 23, 26, 41, 42, 43.

5. Section 15.5: Problems 21, 35, 41, 43, 47.

6. Section 15.7: Problems 9, 11, 13, 14, 15, 19, 21, 31, 39, 43, 45.

7. Section 15.8: Problems 5, 7, 9, 15, 19.

8. Chap 15, additional and advanced exercises: Problems 11, 12.